A car travels at an average speed of 48 miles per hour. How many miles does it travel in 3 hours and 15 minutes?
Found 3 solutions by Alan3354, jim_thompson5910, Maths68:
Answer by Alan3354(69443)

A car travels at an average speed of 48 miles per hour. How many miles does it travel in 3 hours and 15 minutes?
--------------
d = r*t

Answer by jim_thompson5910(35256)

3 hours 15 minutes = 3 + 15/60 = 3 + 0.25 = 3.25 hours

d = rt

d = 48*3.25

d = 156

So it travels 156 miles.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by Maths68(1474)

In three hours it will cover = 3*48 = 144 miles
In 15 minutes or 1/4 of an hour = (1/4)* (48) = 48/4 = 12 miles
Total distance it will cover = 12+144 = 156 miles



OR
3 hours 15 minutes = (60*3)+15 = 180+15 = 195 minutes

In 60 minutes car travels = 48 miles
In 1 minute car will travel = 48/60 = 4/5 miles
In 195 minutes car will travel = (4/5)(195)= (4)(39)= 156 miles